Evaluate the integral "\displaystyle{\int_{0}^{1}te^{t}dt}"


1, 2 ,3 or 4?



1
Expert's answer
2021-09-07T19:15:07-0400

"\int _0^1 te^t dt\\
[ \text{By integration of uv formula}]\\
=[t\int e^t dt] _0^1-\int _0^1 (\frac{dt}{dt}.\int e^t dt) dt\\
=[te^t] _0^1-\int _0^1 e^t dt\\
=e-0-[e^t]_0^1\\
=e-e+1\\
=1"